OpenStudy (anonymous):

what is the y-coordinate of the point at which the line whose equation is 3x-2y-7=0 crosses the y-axis? a.) -7/2 b.) -7/3 c.) 7/3 d.) 7/2 e.) 7

OpenStudy (agent47):

crosses y-axis means x=0, so: 3*0-2y-7=0 -2y=7 y=-7/2